OpenStudy (anonymous):

Solve the differential equation with the initial condition. dy/dx = e^(-x), y(1) = 3

OpenStudy (abb0t):

use separation of variables

OpenStudy (anonymous):

I know the integral of e^(-x) is -e^-x

OpenStudy (abb0t):

\[\int\limits dy = \int\limits \frac{ 1 }{ e^x } dx\]

OpenStudy (anonymous):

plus constant of integration

OpenStudy (anonymous):

yeah, so I am at 3 = -e^(-x) + C

OpenStudy (anonymous):

you could take a natural log of both sides

OpenStudy (anonymous):

sorry, I mean 3 = -e^(-1) + C

OpenStudy (abb0t):

Do what @tomo said. Take the natural log and solve for C. When you solve for C, plug it in.

OpenStudy (anonymous):

so ln(3) = ln(-e^(-1)) + C then solve for C?

OpenStudy (abb0t):

yes

OpenStudy (anonymous):

so I know ln(e^x) is x, but what about ln(-e^(-x))?

OpenStudy (amoodarya):

ln(-e^(-x)) is not valid because e^(-x)>0 so -e^(-x) <0

OpenStudy (anonymous):

Actually ln is not necessary. ok, so dy/dx = e^-x dy = 3^-x*dx y = integral(1/e^x)dx y = -e^-x+c y = c-e^-x y(1) = 3 ok, so we need to figure out c 3 = c-e^-1 3+(1/e) = c

OpenStudy (amoodarya):